OpenStudy (anonymous):

Hello I'm kind of stumbling on this question. What is the slope of the line x = -5? Undefined 0 -5 5 I want to say undefined but I'm not sure

jimthompson5910 (jim_thompson5910):

all points on the line x = -5 have an x coordinate of -5 so points like (-5,0), (-5,1), (-5, 2), etc etc

jimthompson5910 (jim_thompson5910):

a vertical line forms because of this since all vertical lines have undefined slopes, the answer is definitely undefined

jimthompson5910 (jim_thompson5910):

the reason is because slope = rise/run = 1/0 = undefined

OpenStudy (anonymous):

So I was correct ?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

Thanks

jimthompson5910 (jim_thompson5910):

np